# Improving Expressivity of Graph Neural Networks using Localization

### v3

@@ 5.2 Layer k-WL
Consider the $(t+1)^{th}$ round in both the games. $S$ either chooses to play or remove a pebble. If $S$ chooses to remove a pebble from vertex $(v,f(v)),$ then, if $v$ is colored with color $i,$ then $S_{i}$ and $S_{i-1}$ will remove a pebble from vertex $(v,f_{i}(v))$. Again, the position of both the games is same. Now, if $S$ decides to play a pair of pebbles, then each $S_{i}$ also decides to play a pair of pebbles. So, $D$ propose a bijective function and $D_{i}$ proposes a restricted bijective function. Now, suppose $S$ plays a pair of pebbles at $(v_{1},f(v_{1}))$. If $color(v_{1})=i,$ then $S_{i}$ and $S_{i-1}$ also decides to play pebbles at $(v_{1},f_{i}(v_{1}))$. Thus, the position of the game is same as defined. Now, if $S$ wins, then there exists $u$ and $v$ such that either $(u,v)\in E(G)$ and $(f(u),f(v))\notin E(H)$ or $(u,v)\notin E(G)$ and $(f(u),f(v))\in E(H)$. Similarly, there exists $S_{i}$ for which these things happen as the position induced is same. Therefore, $S_{i}$ wins for some $i.$
@@
Thus, from above two lemmas we can say that the expressive power of Layer $k-$WL is the same as local $k-$WL.
@@ 6 Subgraph Counting Algorithms and Characterization of Patterns
Here, we characterize the expressive power of the proposed methods in terms of subgraph as well as induced subgraph counting. In this section, we provide algorithms and characterization of subgraphs that can exactly count the number of patterns appearing as subgraph or induced subgraph. As described above, we can see that the running time is dependent on the size of the local substructure and the value of $k$. The size of the subgraph is dependent on the radius of the patterns. So, we have to take a $r$-hop neighbourhood for each vertex $v$ in the host graph $G$.
In Section 6.1, we show how the value of $k$ can be decided based on the local substructure of the host graph. It is independent of the structure of the pattern. Also, it gives an upper bound on the value of $k$ that can count pattern appearing as subgraph and induced subgraph. In Section 6.2, we first show that doing local count is sufficient for induced subgraph count and later we give the upper bound on $k$ based on the pattern size. Note that the value of $k$ for induced subgraph count is based only on the size of pattern not its structure. In Section 6.3, we again show that locally counting subgraph is sufficient. Also, we explore the value of $k$ based on the structure of pattern. For subgraph counting the structure of pattern can be explored to get a better upperbound for the value of $k$. Later, for the sake of completeness we give algorithm to count triangle, pattern of radius one and $r.$
@@ 6.1 Deciding k based on local substructure of host graph
Here, we explore the local substructure of the host graph in which we are counting pattern appearing as graphs and subgraphs. For a given pattern of radius $r$, we explore $r$-hop neighbourhood around every vertex $v$ in the host graph $G$. If two graphs $G_{1}$ and $G_{2}$ are isomorphic then the number of subgraphs and induced subgraphs of both the graphs are the same. We use the same idea to count number of subgraphs.
Cai et. al. Cai et al. (1992), shown that $\Omega(n)$ dimension is needed to guarantee graph isomorphism. However, for restricted graph classes, we can still guarantee isomorphism for small dimensions. It has been shown that ${3-WL}$ is sufficient for planar graph Kiefer et al. (2019), ${k-WL}$ for graphs with treewidth at most $k$ Kiefer and Neuen (2019), $(3k+4)-WL$ for graphs with rankwidth at most $k$ Grohe and Neuen (2021), and $(4k+3)$ for graphs with Euler genus at most $k$ Grohe and Kiefer (2019). We say these graph classes as good graph classes. Note that, for non-isomorphic graphs, the graphs is not $k-$WL equivalent. Thus, running corresponding $k-$WL can count pattern of radius $r$, appearing as subgraph and induced subgraph.
@@
Consider the subgraph induced by the vertex $v$ and its $r$-hop neighborhood in $G,$ say $G_{v}^{r},$ and the subgraph induced by the vertex $u$ and its $r$-hop neighborhood in $H,$ say $H_{u}^{r}.$ Suppose both structures belong to good graph classes. Now, we run corresponding $k$ based on the local substructure as mentioned in the theorem. If the color histogram of the stable color matches for both the graphs. This implies that both the graphs are isomorphic. Thus, the number of subgraphs and induced subgraphs in both of the substructures are also the same.
Also, we run respective $k-$WL on a colored graph, where vertices at a distance $i$ from $v$ is colored $i$. So, it is at least as expressive as running $k-$WL on an uncolored graph. We can also show that it is strictly more expressive in distinguishing non-isomorphic graphs. Thus, all the $k-$WL mentioned corresponding to good graph classes are sufficient for counting the number of patterns appearing as subgraphs and induced subgraphs. ∎
@@
Running $1-$WL guarantees the exact number of subgraphs and induced subgraphs of all patterns of radius one, when the maximum degree of the host graph is bounded by 5.
@@ 6.2 Counting Induced Subgraphs
The following theorem gives a direct comparison with the $S_{k}$ model Papp and Wattenhofer (2022a), where each node has an attribute which is the count of induced subgraphs of size at most $k.$
@@
Suppose, if possible $G$ and $H$ are Local $k-$WL equivalent and $|P|\leq(k+2),$ where $P$ is the pattern to be counted. We will choose one vertex $v$ as a key vertex. Now, we want to count $P-v$ locally. Assume that the distance from $v$ to any vertex is ${}^{\prime}r^{\prime}$. So, we take the $r$-hop neighborhood of every vertex in $G$ and $H,$ respectively. It is easy to see that the number of induced subgraphs or subgraphs of size $k$ is the same locally if they are $k-$WL equivalent since we do an initial coloring of $k$-tuples based on isomorphism type. Now, suppose $P^{\prime}=P-v$ is of size $(k+1)$.
@@
Now, for each $P^{\prime}$ with $v$ will form $P$ if it has exactly $|N_{P}(v)|$ many vertices of color $1$. Also, as mentioned earlier that $k-$WL is equivalent to $C^{k+1}$ logic, and we have to add unary logic stating that the color of neighbor to be $1$. The $k-$WL and $C^{k+1}$ are equivalent, so we have to add the unary relation putting the condition of the required colors. Again, using Lemma 3 we can say that running $k-$WL locally can count all patterns of size $(k+2)$ appearing as an induced subgraph. ∎
We can see the corollary below in which we mention the set of patterns shown in Chen et al. (2020).
@@
The running time and space requirement for Algorithm 4 is dependent on the value of $k$ and $r$. We can make informed choices for the values of $k$ and $r.$ Notice that the value of $k$ is chosen based on the local substructure. Also, the value of $r$ is the radius of $H.$ Suppose the local substructure is simple (planar, bounded treewidth, bounded rankwidth Theorem 4). In that case, $k-$WL, for small values of $k,$ is sufficient for counting induced subgraph $H.$ Otherwise, we have to run $(|H|-2)-$WL in the worst case.
@@ 6.3 Deciding k based on the pattern for counting subgraphs
where $H^{\prime}$ ranges over all graphs in $H^{\prime}$. This equation has been used to count subgraphs by many authors (Please refer Alon et al. (1997); Curticapean et al. (2017)).
@@
Cai et al. (1992); Dell et al. (2018) For all $k\geq 1$ and for all graphs $G$ and $H$, the following are equivalent:
@@
    Graph $G$ and $H$ are $C^{k+1}$ equivalent***Counting logic with (k+1) variables.
Using Equation 2 and Theorem 7, we arrive at the following theorem:
@@
Suppose $H$ appears as subgraph in $G$. Then, there must exist an injective function matching $u\in V(H)$ to some $v\in V(G)$. Thus, counting locally, we can easily see that every subgraph would be counted. Now to prove Equation 3, it is sufficient to show that for a given subgraph it is counted exactly $|Orbit_{H}(u)|$ many times. Note that two subgraphs are same if and only if their vertex sets and edge sets are same. We fix the vertex set and edge set in $G$, which isomorphic to $H$. Now, consider an automorphism of $H$ which maps $u$ to one of the vertices $u^{\prime}$ in its orbit. Note that we can easily find the updated isomorphic function that maps $u^{\prime}$ to $v$. Now, the number of choices of such $u^{\prime}$ is exactly $|Orbit_{H}(u)|$. Thus, the same subgraph is counted at least $|Orbit_{H}(u)|$ many times. Suppose $x\in V(H)$ is a vertex such that $x\notin Orbit_{H}(u)$. Considering the fixed vertex set and edge set, if we can find an isomorphism, then it is a contradiction to the assumption that $x\notin Orbit_{H}(u)$. Thus, the same subgraph is counted exactly $|Orbit_{H}(u)|$ many times. ∎
Using Theorem 8 and Lemma 5, one can easily see that for counting pattern $H$ as a subgraph, it is sufficient to run Local $k-$WL on the local substructure and count the subgraph locally.
@@
The upper bound on the choice of $k$ for running $k-$WL can be improved from the default $|H|-2$ bound that we used for the induced subgraph count. The value of $k$ is now upper bound by $htw(H)$. Hence, we pick the minimum $k$ based on the local substructure of $G$ as well as the hereditary treewidth of pattern $H$ for computing the subgraph count of $H$ in $G$. The algorithm for counting the subgraph is similar to the induced subgraph.
@@
Local $1-$WL can exactly count the number of patterns $P$ or $P-v$ appearing as a subgraph, when $P$ or $P-v,$ with a dominating vertex $v$, is $K_{1,s}$ and $2K_{2}.$
@@
In a star, $K_{1,s},$ all the leaves are mutually independent. By the definition of homomorphism, the edges are preserved in homomorphic images. So, the only possibility of a homomorphic image of the star is getting another star with less number of leaves. Note that the star is a tree, and its treewidth is one. Also, for $2K_{2}$, the homomorphic image is either itself or the star. So, the treewidth of all its homomorphic images is $1$
@@
We can see that $H-v$ is $K_{1,2}$ when choosing any vertex as a key vertex. Also, the orbit size is $4$. So, we can directly use Lemma 5 to compute the count of the $C_{4}$ in the graph locally and then sum it over all the vertices and divide it by $4$. ∎
@@
For subgraph counting, we can see that for all of the $5$ patterns, there exists a vertex $v$ such that $htw(P-v)=1.$ One can note that the attributed triangle can also be handled using Corollary 1. Since every pattern has a dominating vertex, running $1-$WL on the subgraph induced on the neighborhood is sufficient. Now, we only have to argue for the patterns appearing as induced subgraphs. Note that the induced subgraph count of the triangle and the attributed triangle is same as the subgraph count of the triangle and attributed triangle. ∎
@@
In particular, for counting the number of triangles, we can see that it is enough to count the number of edges in the subgraph induced on the neighbourhood of the vertices. Thus, Local $1-$WL can give the exact count of the number of triangles. For more details, please see 6.4. The running time of $1-$WL depends on the number of iterations, $h$. In general, it takes $O((n+m)\log n)$ time, where $m$ is the number of edges, whereas when we look at it in terms of iteration number it requires $O(nh)$ time.
@@
It requires $O(n)$ time to guarantee the count of patterns which can be written using $2$-variable with a counting quantifier where the depth of the quantifier is constant.
For more details, please see the equivalence between the number of iterations and quantifier depth in Theorem 2. A list of patterns that can be counted as subgraph and induced subgraphs using local $1-$WL and local $2-$WL are mentioned in Table 1. Also, the patterns including 3-star, triangle, chordal $4$-cycle (chordal $C_{4}$), and attributed triangle have been studied in Chen et al. (2020) and have been shown that it cannot be counted by $1-$WL.
@@
Restriction on $G_{v}^{r}$ k Patterns, $H$
Induced
subgraph
Subgraph Reference
$G_{v}^{r}$ is amenable 1
All
Corollary 1
Max degree $\leq$ 5 1
Patterns with a dominating vertex
Corollary 2
Max degree $\leq$ 15 2
Pattern with a dominating vertex
Corollary 2
No restriction 2
3-star, triangle, tailed triangle, chordal cycle,
attributed triangle
Corollary 3
No restriction 1
Either $H$ or $H-v$ is $K_{1,s}$ or $2K_{2},$ where $v$ is the
dominating vertex
Corollary 4
@@
No restriction 1
3-star, tailed triangle, chordal cycle
Corollary 6
No restriction 1
triangle, attributed triangle
Corollary 6

### v5

@@ 5.2 Layer k-WL
Consider the $(t+1)^{th}$ round in both the games. $S$ either chooses to play or remove a pebble. If $S$ chooses to remove a pebble from vertex $(v,f(v)),$ then, if $v$ is colored with color $i,$ then $S_{i}$ and $S_{i-1}$ will remove a pebble from vertex $(v,f_{i}(v))$. Again, the position of both the games is the same. Now, if $S$ decides to play a pair of pebbles, then each $S_{i}$ also decides to play a pair of pebbles. So, $D$ propose a bijective function, and $D_{i}$ proposes a restricted bijective function. Now, suppose $S$ plays a pair of pebbles at $(v_{1},f(v_{1}))$. If $color(v_{1})=i,$ then $S_{i}$ and $S_{i-1}$ also decides to play pebbles at $(v_{1},f_{i}(v_{1}))$. Thus, the position of the game is the same as defined. Now, if $S$ wins, then there exists $u$ and $v$ such that either $(u,v)\in E(G)$ and $(f(u),f(v))\notin E(H)$ or $(u,v)\notin E(G)$ and $(f(u),f(v))\in E(H)$. Similarly, there exists $S_{i}$ for which these things happen as the position induced is the same. Therefore, $S_{i}$ wins for some $i.$
@@
Thus, from the above two lemmas, we can say that the expressive power of Layer $k$-WL is the same as local $k$-WL.
@@ 6 Subgraph Counting Algorithms and Characterization of Patterns
Here, we characterize the expressive power of the proposed methods in terms of subgraph as well as induced subgraph counting. In this section, we provide algorithms and characterization of subgraphs that can exactly count the number of patterns appearing as subgraphs or induced subgraphs. As described above, we can see that the running time is dependent on the size of the local substructure and the value of $k$. The size of the subgraph is dependent on the radius of the patterns. So, we have to take a $r$-hop neighbourhood for each vertex $v$ in the host graph $G$.
In 6.1, we show how the value of $k$ can be decided based on the local substructure of the host graph. It is independent of the structure of the pattern. Also, it gives an upper bound on the value of $k$ that can count patterns appearing as subgraphs and induced subgraphs. In 6.2, we first show that doing local count is sufficient for induced subgraph count, and later, we give the upper bound on $k$ based on the pattern size. Note that the value of $k$ for induced subgraph count is based only on the size of the pattern, not its structure. In 6.3, we again show that locally counting subgraph is sufficient. Also, we explore the value of $k$ based on the structure of the pattern. For subgraph counting, the structure of the pattern can be explored to get a better upper bound for the value of $k$. Later, for the sake of completeness, we give an algorithm to count triangles, the pattern of radius one and $r.$
@@ 6.1 Deciding k based on local substructure of host graph
Here, we explore the local substructure of the host graph in which we are counting patterns appearing as graphs and subgraphs. For a given pattern of radius $r$, we explore the $r$-hop neighbourhood around every vertex $v$ in the host graph $G$. If two graphs $G_{1}$ and $G_{2}$ are isomorphic, then the number of subgraphs and induced subgraphs of both the graphs are the same. We use the same idea to count the number of subgraphs.
cai1992optimal, shown that $\Omega(n)$ dimension is needed to guarantee graph isomorphism. However, for restricted graph classes, we can still guarantee isomorphism for small dimensions. It has been shown that 3WL{3-WL} is sufficient for planar graph (kiefer2019weisfeiler), kWL{k-WL} for graphs with treewidth at most $k$ (decompose_graphs), (3k+4)WL(3k+4)-WL for graphs with rankwidth at most $k$ (rankwidth_WL), and $(4k+3)$ for graphs with Euler genus at most $k$ (grohe2019linear). We say these graph classes as good graph classes. Note that, for non-isomorphic graphs, the graphs is not $k$-WL equivalent. Thus, running corresponding $k$-WL can count the pattern of radius $r$, appearing as a subgraph and induced subgraph.
@@
Consider the subgraph induced by the vertex $v$ and its $r$-hop neighbourhood in $G,$ say $G_{v}^{r},$ and the subgraph induced by the vertex $u$ and its $r$-hop neighbourhood in $H,$ say $H_{u}^{r}.$ Suppose both structures belong to good graph classes. Now, we run corresponding $k$ based on the local substructure as mentioned in the theorem. If the colour histogram of the stable colour matches for both graphs. This implies that both the graphs are isomorphic. Thus, the number of subgraphs and induced subgraphs in both of the substructures are also the same.
Also, we run respective $k$-WL on a coloured graph, where vertices at a distance $i$ from $v$ are coloured $i$. So, it is at least as expressive as running $k$-WL on an uncoloured graph. We can also show that it is strictly more expressive in distinguishing non-isomorphic graphs. Thus, all the $k$-WL mentioned corresponding to good graph classes are sufficient for counting the number of patterns appearing as subgraphs and induced subgraphs. ∎
@@
Running 11-WL guarantees the exact number of subgraphs and induced subgraphs of all patterns of radius one, when the maximum degree of the host graph is bounded by 5.
@@ 6.2 Counting Induced Subgraphs
The following theorem gives a direct comparison with the $S_{k}$ model (papp2022theory), where each node has an attribute which is the count of induced subgraphs of size at most $k.$
@@
Suppose, if possible $G$ and $H$ are Local $k$-WL equivalent and $|P|\leq(k+2),$ where $P$ is the pattern to be counted. We will choose one vertex $v$ as a key vertex. Now, we want to count PvP-v locally. Assume that the distance from $v$ to any vertex is $^{\prime}r^{\prime}$. So, we take the $r$-hop neighborhood of every vertex in $G$ and $H,$ respectively. It is easy to see that the number of induced subgraphs or subgraphs of size $k$ is the same locally if they are $k$-WL equivalent since we do an initial coloring of $k$-tuples based on isomorphism type. Now, suppose $P^{\prime}=P-v$ is of size $(k+1)$.
@@
Now, for each $P^{\prime}$ with $v$ will form $P$ if it has exactly $|N_{P}(v)|$ many vertices of color $1$. Also, as mentioned earlier that $k$-WL is equivalent to $C^{k+1}$ logic, and we have to add unary logic stating that the color of neighbor to be $1$. The $k$-WL and $C^{k+1}$ are equivalent, so we have to add the unary relation putting the condition of the required colors. Again, using 3 we can say that running $k$-WL locally can count all patterns of size $(k+2)$ appearing as an induced subgraph. ∎
We can see the corollary below in which we mention the set of patterns shown in (chen2020can).
@@
The running time and space requirement for Algorithm 4 is dependent on the value of $k$ and $r$. We can make informed choices for the values of $k$ and $r.$ Notice that the value of $k$ is chosen based on the local substructure. Also, the value of $r$ is the radius of $H.$ Suppose the local substructure is simple (planar, bounded treewidth, bounded rankwidth 4). In that case, $k$-WL, for small values of $k,$ is sufficient for counting induced subgraph $H.$ Otherwise, we have to run (|H|2)(|H|-2)-WL in the worst case.
@@ 6.3 Deciding k based on the pattern for counting subgraphs
where $H^{\prime}$ ranges over all graphs in $H^{\prime}$. This equation has been used to count subgraphs by many authors (Please refer (Alon1997; curticapean2017homomorphisms)).
@@
(cai1992optimal; Dell2018LovszMW) For all $k\geq 1$ and for all graphs $G$ and $H$, the following are equivalent:
@@
    Graph $G$ and $H$ are $C^{k+1}$ equivalent [^2] Counting logic with (k+1) variables.
Using 3 and 7, we arrive at the following theorem:
@@
Suppose $H$ appears as subgraph in $G$. Then, there must exist an injective function matching $u\in V(H)$ to some $v\in V(G)$. Thus, counting locally, we can easily see that every subgraph would be counted. Now to prove 4, it is sufficient to show that for a given subgraph it is counted exactly $|Orbit_{H}(u)|$ many times. Note that two subgraphs are same if and only if their vertex sets and edge sets are same. We fix the vertex set and edge set in $G$, which isomorphic to $H$. Now, consider an automorphism of $H$ which maps $u$ to one of the vertices $u^{\prime}$ in its orbit. Note that we can easily find the updated isomorphic function that maps $u^{\prime}$ to $v$. Now, the number of choices of such $u^{\prime}$ is exactly $|Orbit_{H}(u)|$. Thus, the same subgraph is counted at least $|Orbit_{H}(u)|$ many times. Suppose $x\in V(H)$ is a vertex such that $x\notin Orbit_{H}(u)$. Considering the fixed vertex set and edge set, if we can find an isomorphism, then it is a contradiction to the assumption that $x\notin Orbit_{H}(u)$. Thus, the same subgraph is counted exactly $|Orbit_{H}(u)|$ many times. ∎
Using 8 and 5, one can easily see that for counting pattern $H$ as a subgraph, it is sufficient to run Local $k$-WL on the local substructure and count the subgraph locally.
@@
The upper bound on the choice of $k$ for running $k$-WL can be improved from the default |H|2|H|-2 bound that we used for the induced subgraph count. The value of $k$ is now upper bound by $htw(H)$. Hence, we pick the minimum $k$ based on the local substructure of $G$ as well as the hereditary treewidth of pattern $H$ for computing the subgraph count of $H$ in $G$. The algorithm for counting the subgraph is similar to the induced subgraph.
@@
Local $1$-WL can exactly count the number of patterns $P$ or PvP-v appearing as a subgraph, when $P$ or Pv,P-v, with a dominating vertex $v$, is $K_{1,s}$ and $2K_{2}.$
@@
In a star, $K_{1,s},$ all the leaves are mutually independent. By the definition of homomorphism, the edges are preserved in homomorphic images. So, the only possibility of a homomorphic image of the star is getting another star with less number of leaves. Note that the star is a tree, and its treewidth is one. Also, for $2K_{2}$, the homomorphic image is either itself or the star. So, the treewidth of all its homomorphic images is $1.$
@@
We can see that HvH-v is $K_{1,2}$ when choosing any vertex as a key vertex. Also, the orbit size is $4$. So, we can directly use Lemma 5 to compute the count of the $C_{4}$ in the graph locally and then sum it over all the vertices and divide it by $4$. ∎
@@
For subgraph counting, we can see that for all of the $5$ patterns, there exists a vertex $v$ such that htw(Pv)=1.htw(P-v)=1. One can note that the attributed triangle can also be handled using Corollary 1. Since every pattern has a dominating vertex, running $1$-WL on the subgraph induced on the neighbourhood is sufficient. Now, we only have to argue for the patterns appearing as induced subgraphs. Note that the induced subgraph count of the triangle and the attributed triangle is same as the subgraph count of the triangle and attributed triangle. ∎
@@
In particular, for counting the number of triangles, we can see that it is enough to count the number of edges in the subgraph induced on the neighbourhood of the vertices. Thus, Local $1$-WL can give the exact count of the number of triangles. For more details, please see 6.4.1. The running time of $1$-WL depends on the number of iterations, $h$. In general, it takes $O((n+m)\log n)$ time, where $m$ is the number of edges, whereas when we look at it in terms of iteration number it requires $O(nh)$ time.
@@
It requires $O(n)$ time to guarantee the count of patterns which can be written using 22-variable with a counting quantifier where the depth of the quantifier is constant.
For more details, please see the equivalence between the number of iterations and quantifier depth in 2. A list of patterns that can be counted as subgraph and induced subgraphs using local $1$-WL and local $2$-WL are mentioned in 1. Also, the patterns including 3-star, triangle, chordal 44-cycle (chordal $C_{4}$), and attributed triangle have been studied in (chen2020can) and have been shown that it cannot be counted by 11-WL.
@@
Restriction on $G_{v}^{r}$ k Patterns, $H$ Induced subgraph Subgraph Reference
$G_{v}^{r}$ is amenable 1 All Corollary 1
Max degree $\leq 5$ 1 Patterns with a dominating vertex Corollary 2
Max degree $\leq 15$ 2 Pattern with a dominating vertex Corollary 2
No restriction 2 3-star, triangle, tailed triangle, chordal cycle, attributed triangle Corollary 3
No restriction 1 Either $H$ or HvH-v is $K_{1,s}$ or $2K_{2},$ where $v$ is the dominating vertex Corollary 4
@@
No restriction 1 3-star, tailed triangle, chordal cycle Corollary 6
No restriction 1 triangle, attributed triangle Corollary 6